If you have 2.75 moles of Sn(ClO3)2, what is its mass?
Answer #1

Sn(ClO3)2

1 mole of above compound equals to molecular mass

molecular mass=118.71 +2*35.45 +6*16=285.61 g/mol

1mol \rightarrow 285.61 g

2.75 moles of Sn(ClO3)2 has mass equals to 285.61*2.75=785.427g